OpenStudy (anonymous):

find the indicated outputs for f(x)=5²-5x f(0)= f(-1)= f(2)=

OpenStudy (anonymous):

f(0)=5²-5*0=5³=25 f(-1)=25-5*-1=25+5=30 f(2)=25-5*x=25-10=15

OpenStudy (anonymous):

Thank you!